A distance runner ran 188 miles in 24 hours. After 8 hours he was approximately twice as far from the finish line as he was from the start. How far did he run?
I have problems with word problems along with just learning algebra.

Answer by stanbon(75887)

A distance runner ran 188 miles in 24 hours.
---
rate = 188/24 = 7.83333 mph
--------
After 8 hours he was approximately twice as far from the finish line as he was from the start.
Comment: That means he has run 1/3 of the total and still has 2/3 of the
total distance to run.
--------------------------
How far did he run?
-------
Let x be the total distance of his run from start to finish.
distance = distance
8*7.83333 = (1/3)x
He has run 8*7.8333333 = 62 2/3 miles
=======================================
Cheers,
Stan H.
=================
------------------